OpenStudy (anonymous):

What is the value of a in the equation 2 cos a = 1?

OpenStudy (anonymous):

a = pi / 3 or a = 5pi / 3

OpenStudy (anonymous):

lazbumtac how did you get that??

OpenStudy (anonymous):

Haha I have the unit circle memorized in my head

OpenStudy (anonymous):

If you look at a unit circle, then cos(pi /3) = 1/2. Then (1/2) * 2 = 1. cos(5pi /3) is also = 1/2

OpenStudy (anonymous):

thank you!